OpenStudy (anonymous):

-(10)^-1 , simple problem, i know. just wanna make sure im on the right track

OpenStudy (anonymous):

what did you get?

OpenStudy (okra77):

anything to a negative exponent can be made positive by moving to the denominator 1/-10 

OpenStudy (smileyxl3):

I get \[-\frac{ 1 }{10 }\]

OpenStudy (anonymous):

the negative stays because it's not in the parentheses \[-(10)^{-1} = -\frac{1}{10^{1}} = -\frac{1}{10}\]